# Bayesian Ensembling: Insights from Online Optimization and Empirical Bayes

### v1

@@ Abstract
We revisit the classical problem of Bayesian ensembles and address the challenge of learning optimal combinations of Bayesian models in an online, continual learning setting. To this end, we reinterpret existing approaches such as Bayesian model averaging (BMA) and Bayesian stacking through a novel empirical Bayes lens, shedding new light on the limitations and pathologies of BMA. Further motivated by insights from online optimization, we propose Online Bayesian Stacking (OBS), a method that optimizes the log-score over predictive distributions to adaptively combine Bayesian models. A key contribution of our work is establishing a novel connection between OBS and portfolio selection, bridging Bayesian ensemble learning with a rich, well-studied theoretical framework that offers efficient algorithms and extensive regret analysis. We further clarify the relationship between OBS and online BMA, showing that they optimize related but distinct cost functions. Through theoretical analysis and empirical evaluation, we identify scenarios where OBS outperforms online BMA and provide principled guidance on when practitioners should prefer one approach over the other.
@@ 1 Introduction
Combining the opinions of multiple models is a pervasive problem in the statistical sciences, with many different names, approaches, and applications. In signal processing, for example, a commonly encountered problem is one of sensor fusion, where information reported from several different sensors must be combined to obtain the best possible estimate (Khaleghi et al., 2013). In econometrics, the problem is often known as forecast combination, spurred by the seminal work of Bates & Granger (1969). The Bayesian literature often refers to the problem as opinion pooling, and received attention even in the early days of Bayesian statistics (Stone, 1961; DeGroot, 1974; Genest & Zidek, 1986). While many of the above approaches originated in the combination of point estimates, the combination of probability distributions has received significant attention since (Koliander et al., 2022).
In machine learning, this issue is commonly known as ensembling, with much recent interest due to the development of diverse models, architectures, and training modalities (Dietterich, 2000). Many ensembling approaches exist, hinging on different assumptions and asymptotic guarantees. We will restrict our attention to Bayesian machine learning, where the goal is to combine the estimates of $K$ different probabilistic models $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$. The classical approach in this setting, known as Bayesian model averaging (BMA), is to weigh the estimates of each model according to their marginal likelihood (Hoeting et al., 1999). When the data were generated by one of the models $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$, this is the “correct” way to combine models from a Bayesian perspective, and is optimal, in the sense of choosing the correct model with probability $1$ in the limit of infinite data (Yao et al., 2018).
@@
In this paper, we further study the online stacking problem, introducing a novel analysis via empirical Bayes, before moving to a new analysis of Bayesian stacking in the online setting. We summarize our contributions as follows:
@@
    We show that Online Bayesian Stacking (OBS) corresponds exactly to the well-studied problem of (universal) online portfolio selection (OPS). This connection allows us to reinterpret Bayesian ensemble learning through the lens of online convex optimization and to leverage efficient, theoretically grounded algorithms (e.g., Exponentiated Gradient and the Online Newton Step). Building on this connection, we discuss how regret bounds from the OPS literature can be applied to OBS.
@@
    We introduce a simple yet compelling argument via empirical Bayes to explain why BMA collapses, and how Bayesian stacking aids in avoiding this common problem.
@@
    We perform an extensive empirical analysis of OBS using ensembles of state-of-the-art models, including Gaussian processes, variational Bayesian neural networks, and stochastic volatility forecasting models estimated with sequential Monte Carlo. We show that, in all cases tested, OBS significantly outperforms online BMA (O-BMA) and dynamic model averaging (DMA), with negligible additional computational cost. We also show that OBS can be particularly beneficial in non-stationary environments, often outperforming any set of fixed weights.
The rest of this article is structured as follows: in Section 2, we review the problem of Bayesian ensembling, discussing the BMA and stacking approaches. In Section 3, we show how the online variant of Bayesian stacking corresponds to the well-known (universal, online) portfolio selection problem from online convex optimization, and discuss corresponding insights that show OBS’s relation to O-BMA. In Section 4, we exploit this connection to derive a novel comparison of BMA and Bayesian stacking from an empirical Bayes perspective. In Section 5, we carry out a number of realistic experiments on synthesized and real data, before finally concluding in Section 6.
@@ 2 Bayesian Ensembles: A Review of Existing Methods
In this section, we review the ideas of BMA and Bayesian stacking Yao et al. (2018), and discuss some related work.
@@ 2.1 Bayesian Model Averaging
Bayesian ensembling methods combine a set of Bayesian models for predictive inference. We will focus on methods that create linear mixtures of posterior predictive distributions: given models $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$ each mapping $\boldsymbol{\mathbf{x}}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ to a probability (density) over $y\in\mathcal{Y}\subseteq\mathbb{R}^{r}$, trained on a dataset $\mathcal{D}$, the task of Bayesian ensembling is thus to find a weight vector $\mathbf{w}$ in the $(K-1)$-dimensional simplex $\mathbb{S}^{K}$, which induces a posterior predictive distribution
@@
where $p_{k}(\mathcal{D})\coloneq p(\mathcal{D}|\mathcal{M}_{k})$ is the evidence of ${\cal M}_{k}$ and $\Pr(\mathcal{M}_{k})$ is the prior probability of ${\cal M}_{k}$.
The appeal of BMA to a Bayesian is immediate: if $\Pr(\mathcal{M})$ has support over the data-generating distributions, then the resulting mixture Equation 1 is a straightforward application of Bayes’ rule, and thus optimal in an information-theoretic sense (Zellner, 1988).
Much of our work focuses on the online setting, where BMA has several additional advantages. In this setting, we first obtain a new data point $\mathbf{x}_{t+1}$, with which we must make a prediction using the available data $\mathcal{D}_{t}$. After making a prediction, the value $y_{t+1}$ is revealed and the dataset $\mathcal{D}_{t+1}\coloneq\mathcal{D}\cup\{\mathbf{x}_{t+1},y_{t+1}\}$ is updated. In this case, can compute $w_{k}^{\text{BMA}}$ recursively using posterior predictive distributions, which allows us to apply exact BMA at time $t$ using the currently-available information: denoting the weights at time $t$ as $w_{t,k}$, with $w_{0,k}=\Pr(\mathcal{M}_{k})$, (2) becomes
@@
Two main problems arise when using BMA: first, estimation via Equation 2 or Equation 3 requires access to predictive distributions, either through the marginal likelihood (i.e., the prior predictive) or the posterior predictive. This is often a surmountable problem, as marginal likelihoods are usually available in conjugate models, and BMA with approximated marginal likelihoods has also performed well empirically (Gómez-Rubio et al., 2020).
Perhaps more importantly, BMA is only optimal in terms of predictive error in the so-called “M-closed” setting, where the data were generated by one of the models $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$ (Yao et al., 2018; Minka, 2000). Indeed, in the limit of infinite data, BMA weights concentrate on the single model that most closely resembles the data-generating process. We provide a novel empirical Bayes-based analysis of this fact in Section 4. This may result in arbitrarily poor posterior predictive accuracy of an ensemble using $\mathbf{w}^{\text{BMA}}$ relative to some other set of weights $\mathbf{w}^{*}$. Additionally, O-BMA can numerically collapse to the “wrong” model, never recovering due to numerical underflow (Waxman & Djurić, 2024).
@@ 2.2 Bayesian Stacking
Bayesian stacking (Yao et al., 2018) (similarly explored by (Clydec & Iversen, 2013; Le & Clarke, 2017)) presents an alternative way to derive a weight vector $\mathbf{w}$. In particular, Bayesian stacking finds the optimal weight vector $\boldsymbol{\mathbf{w}}^{*}$ in the $(K-1)$-dimensional simplex $\mathbb{S}^{K}$ by maximizing some score $\mathcal{S}(\boldsymbol{\mathbf{w}},\mathcal{D}^{\prime})$ over a separate dataset $\mathcal{D}^{\prime}$:
@@
Particular attention is given to the log-score; for the predictive dataset $\mathcal{D}^{\prime}=\{(\boldsymbol{\mathbf{x}}_{n},y_{n})\}_{n=1}^{N}$ [^1] We will typically use $n=1,\dots,N$ to denote datasets that are not processed sequentially/online, and $t=1,\dots,T$ for their sequential/online counterparts., the corresponding optimization problem is
@@
where $p_{k}(y_{n}\,|\,\boldsymbol{\mathbf{x}}_{n},\mathcal{D})$ is the predictive distribution of model $\mathcal{M}_{k}$. This can be recognized as finding the mixture of estimators with the highest predictive likelihood over $\mathcal{D}^{\prime}$. This in turn minimizes the KL divergence between the mixture $q(y\,|\,\boldsymbol{\mathbf{x}})=\sum_{k}w_{k}p_{k}(y\,|\,\boldsymbol{\mathbf{x}},\mathcal{D})$ and the predictive distribution, $p(y\,|\,\boldsymbol{\mathbf{x}})$, which represents the (unknown) true generating mechanism of $y$ (Yao et al., 2018).
The optimization problem (5) is convex and provides nice asymptotic guarantees, but “wastes” data by requiring two separate datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. Yao et al. (2018) address this by showing the score in (5) is well-approximated by the leave-one-out (LOO) predictive density $p(y_{n}\,|\,\boldsymbol{\mathbf{x}}_{n},\mathcal{D}_{-n})$, where $\mathcal{D}_{-n}=\mathcal{D}\setminus\{(\boldsymbol{\mathbf{x}}_{n},y_{n})\}$. For many models, the LOO predictive density can be efficiently estimated with Pareto-smoothed importance sampling (Vehtari et al., 2024), resulting in an efficient optimization problem with a single dataset $\mathcal{D}$. The resulting mixture of estimators has empirically shown performance superior to other combination methods, leading to several recent applications.
@@ 2.3 Bayesian Stacking for Time Series & Online Bayesian Stacking
In a contributed discussion to Yao et al. (2018), Ferreira (2018) discusses Bayesian stacking for time series, where forecasting densities naturally serve as predictive distributions, i.e., (5) becomes
@@
where the first $t^{*}$ data are devoted to (pre-)training each model and $\mathcal{D}_{1:t-1}$ denotes the first $t-1$ data points. This has notable similarities to pooling methods described by Hall & Mitchell (2007); Geweke & Amisano (2011), with the exception of the summation starting with index $t=1$, reminiscent of the LOO approach. This method was applied by Geweke & Amisano (2012), where optimal weights were computed quarterly. “Windowed” approaches for autocorrelated time series data where only the last $T_{p}$ points are considered have also been deployed (Jore et al., 2010; Aastveit et al., 2014), but these are suboptimal in the more general case where data may be exchangeable conditioned on $\boldsymbol{\mathbf{x}}_{t}$. Related to windowed approaches is dynamic model averaging, which is O-BMA with forgetting factors (Raftery et al., 2010). To the best of our knowledge, Bayesian stacking and its time series variants have not been applied in the online setting, where weights are estimated as new data become available. We will refer to approaches to the online problem as online Bayesian stacking (OBS).
As we will see, OBS is a special case of online convex optimization (OCO) (Hazan, 2022), where “learning from experts” is well-studied. We show the popular Hedge algorithm (Freund & Schapire, 1997) generalizes BMA with a learning rate, but optimizes a different loss function from OBS. Our proposed OBS is differentiated from existing approaches in several ways. First, it directly emulates Bayesian stacking, of recent interest to the Bayesian community. Second, our connection to OCO yields extremely efficient implementation, unlike methods based on data-driven portfolio selection (e.g., Baştürk et al. (2019)) that rely on particle filtering. Finally, our approach is more general for online learning, whereas data-driven strategies rely on the autocorrelation of time series data.
@@ 3 Insights from Online Convex Optimization
As Bayesian stacking is natively framed as an optimization problem it is natural to study OBS from the optimization perspective as well. It turns out that our studies here are fruitful: by interpreting posterior predictive values as asset prices, OBS corresponds exactly to a classical problem of portfolio selection. An OCO perspective additionally rediscovers a development of Vovk (2001), which shows BMA is exactly the Hedge algorithm with a specific choice in learning rate.
@@ 3.1 The Portfolio Selection Problem
We will first review the basic ideas of online portfolio selection (OPS), but our discussion will be necessarily brief; the interested reader is pointed to the recent survey of Li & Hoi (2014), or the excellent lecture notes of Hazan (2022); Orabona (2019).
Problem Statement Consider an investment manager overseeing $K$ stocks, aiming to maximize total future wealth. They create the best constantly rebalanced portfolio (BCRP), i.e., at each time step they re-allocate current wealth such that $w_{k}$ is in stock $k$ (classic portfolio selection assumes no transaction fees). Assets are allocated based on price-relative vectors $\boldsymbol{\mathbf{r}}_{t}$, where $r_{t,k}$ is stock $k$’s relative value increase from day $t-1$ to $t$. The goal is to maximize the (multiplicative) wealth at time $T$, or more conveniently, the (additive) log-wealth:
@@
In the online version of the portfolio selection problem, we seek to find $\boldsymbol{\mathbf{w}}_{t}$ which minimizes the regret with respect to the optimal weights $\boldsymbol{\mathbf{w}}^{*}$,
@@
Although many approaches explicitly model returns as stochastic processes, others allow $r_{t,k}$ to be chosen arbitrarily, and indeed by an adversary. Algorithms that obtain sublinear regret with respect to any non-negative sequence of returns $r_{t,k}$ are known as universal (Orabona, 2019, p. 144). We draw an analogy between the price-relative vectors $\boldsymbol{\mathbf{r}}_{t}$ and the predictive likelihoods $p(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})$, which makes the adversarial case of particular interest to us.
Algorithms The portfolio selection problem has an attractive structure from the optimization perspective: the log wealth is a concave function defined over the simplex, a convex set. OPS therefore falls under the (online) convex optimization (OCO) umbrella (Hazan, 2022; Orabona, 2019). To keep with standard terminology, we will introduce each method as minimizing the convex loss function given by $-\log\text{wealth}_{T}$, equivalent to maximizing the concave function $\log\text{wealth}_{T}$.
The simplest OCO algorithm is online (sub)gradient descent (OGD), a straightforward extension of the classical gradient descent algorithm to a sequence of losses $\ell_{1},\dots,\ell_{T}$, each a function of $\boldsymbol{\mathbf{x}}$ belonging to a convex set $\mathcal{K}$. In our case, $\boldsymbol{\mathbf{x}}_{t}$ represents weights $\boldsymbol{\mathbf{w}}_{t}$, $\ell_{t}$ is the negative log-return, $-\log(\boldsymbol{\mathbf{w}}_{t}^{\top}\boldsymbol{\mathbf{r}}_{t})$, and $\mathcal{K}$ is the simplex $\mathbb{S}^{K}$. At each time instant, $\boldsymbol{\mathbf{x}}_{t}$ is updated in the direction of the gradient $\nabla\ell_{t}$ and then projected back onto the convex set $\mathcal{K}$ (e.g., (Hazan, 2022, Algorithm 2.3)).
In many problems, it turns out beneficial to consider regularizers that account for the geometry of $\mathcal{K}$ explicitly, resulting in different update steps. This is an online form of the classical mirror descent (Nemirovsky et al., 1983). When $\mathcal{K}$ is the simplex $\mathbb{S}^{K}$ (like in OPS), one popular choice is entropic regularization (e.g., (Hazan, 2022, Section 5.4.2)). The resulting algorithm is known as Exponentiated Gradients (EG) (Helmbold et al., 1998), which uses the update step
@@
where $\odot$ is the elementwise product and $\eta$ is a learning rate. For the cost function $\ell_{t}(\boldsymbol{\mathbf{w}})=-\log\boldsymbol{\mathbf{w}}_{t}^{\top}\boldsymbol{\mathbf{r}}_{t}$, we have $\nabla_{\boldsymbol{\mathbf{w}}}\ell_{t}=-\boldsymbol{\mathbf{r}}_{t}/\boldsymbol{\mathbf{w}}_{t}^{\top}\boldsymbol{\mathbf{r}}_{t}$, and therefore,
@@
The EG algorithm is conceptually simple with relatively good regret bounds whenever the gradient is bounded. However, more modern algorithms such as the online Newton step (ONS) (Hazan et al., 2007) can provide tighter bounds on the regret.

### v2

@@ Abstract
We revisit the classical problem of Bayesian ensembles and address the challenge of learning optimal combinations of Bayesian models in an online learning setting. To this end, we reinterpret existing approaches such as Bayesian model averaging (BMA) and Bayesian stacking through a novel empirical Bayes lens, shedding new light on the limitations and pathologies of BMA. Further motivated by insights from online optimization, we propose Online Bayesian Stacking (OBS), a method that optimizes the log-score over predictive distributions to adaptively combine Bayesian models. A key contribution of our work is establishing a novel connection between OBS and portfolio selection, bridging Bayesian ensemble learning with a rich, well-studied theoretical framework that offers efficient algorithms and extensive regret analysis. We further clarify the relationship between OBS and online BMA, showing that they optimize related but distinct cost functions. Through theoretical analysis and empirical evaluation, we identify scenarios where OBS outperforms online BMA and provide principled methods and guidance on when practitioners should prefer one approach over the other.
@@ 1 Introduction
Combining the opinions of multiple models is a pervasive problem in the statistical sciences, with many different names, approaches, and applications. In signal processing, for example, a commonly encountered problem is one of sensor fusion, where information reported from several different sensors must be combined to obtain the best possible estimate (Khaleghi et al., 2013). In econometrics, the problem is often known as forecast combination, spurred by the seminal work of Bates & Granger (1969). The Bayesian literature often refers to the problem as opinion pooling, whichreceived attention even in the early days of Bayesian statistics (Stone, 1961; DeGroot, 1974; Genest & Zidek, 1986). While many of the above approaches originated in the combination of point estimates, the combination of probability distributions has received significant attention since (Koliander et al., 2022).
In machine learning, this issue is commonly known as ensembling, with much recent interest due to the development of diverse models, architectures, and training modalities (Dietterich, 2000). Many ensembling approaches exist, hinging on different assumptions and asymptotic guarantees. We will restrict our attention to Bayesian machine learning, where the goal is to combine the estimates of $K$ different probabilistic models $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$. The classical approach in this setting, known as Bayesian model averaging (BMA), is to weigh the estimates of each model according to their marginal likelihood (Hoeting et al., 1999). When the data were generated by one of the models $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$, this is the “correct” way to combine models from a Bayesian perspective and is optimal, in the sense of choosing the correct model with probability $1$ in the limit of infinite data (Yao et al., 2018).
@@
In this paper, we further study the online stacking problem, introducing a novel analysis via empirical Bayes before moving to a new analysis of Bayesian stacking in the online setting. We summarize our contributions as follows:
@@
    We show that Online Bayesian Stacking (OBS) corresponds exactly to the well-studied problem of (universal) online portfolio selection (OPS). This connection allows us to reinterpret Bayesian ensemble learning through the lens of online convex optimization and to leverage efficient, theoretically grounded algorithms (e.g., Exponentiated Gradient and the Online Newton Step).
@@
    We show theoretically how one can choose between OBS and online BMA with constant regret. Building on this connection, we discuss how regret bounds from the OPS literature can be applied to OBS.
@@
    We introduce a simple yet compelling argument via empirical Bayes to explain why BMA collapses and how Bayesian stacking aids in avoiding this common problem.
  4. 4.
    We perform an extensive empirical analysis of OBS using ensembles of state-of-the-art models, including Gaussian processes, variational Bayesian neural networks, and stochastic volatility forecasting models estimated with sequential Monte Carlo. We show that, in all cases tested, OBS significantly outperforms online BMA (O-BMA) and dynamic model averaging (DMA), with additional computational cost that is often negligible in comparison to online training/prediction of ML models. We also show that OBS can be particularly beneficial in non-stationary environments, often outperforming any set of fixed weights.
The rest of this article is structured as follows: in Section˜2, we review the problem of Bayesian ensembling, discussing the BMA and stacking approaches. In Section˜3, we show how the online variant of Bayesian stacking corresponds to the well-known (universal, online) portfolio selection problem from online convex optimization, and discuss corresponding insights that show OBS’s relation to O-BMA. In Section˜4, we exploit this connection to derive a novel comparison of BMA and Bayesian stacking from an empirical Bayes perspective. In Section˜5, we carry out a number of realistic experiments on synthesized and real data before finally concluding in Section˜6.
@@ 2 Bayesian Ensembles: A Review of Existing Methods
In this section, we review the ideas of BMA and Bayesian stacking Yao et al. (2018) and discuss some related work.
@@ 2.1 Bayesian Model Averaging
Bayesian ensembling methods combine a set of Bayesian models for predictive inference. We will focus on methods that create linear mixtures of posterior predictive distributions: given models $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$ each mapping $\boldsymbol{\mathbf{x}}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ to a probability (density) over $y\in\mathcal{Y}\subseteq\mathbb{R}^{r}$, trained on a dataset $\mathcal{D}$, the task of Bayesian ensembling is thus to find a weight vector $\mathbf{w}$ in the (K1)(K-1)-dimensional simplex $\mathbb{S}^{K}$, which induces a posterior predictive distribution
@@
where $p_{k}(\mathcal{D})\coloneq p(\mathcal{D}|\mathcal{M}_{k})$ is the evidence of ${\cal M}_{k}$, and $\Pr(\mathcal{M}_{k})$ is the prior probability of ${\cal M}_{k}$.
The appeal of BMA to a Bayesian is immediate: if $\Pr(\mathcal{M})$ has support over the data-generating distributions, then the resulting mixture Equation˜1 is a straightforward application of Bayes’ rule and is thus optimal in an information-theoretic sense (Zellner, 1988).
Much of our work focuses on the online setting, where BMA has several additional advantages. In this setting, we first obtain a new data point $\mathbf{x}_{t+1}$, with which we must make a prediction using the available data $\mathcal{D}_{t}$. After making a prediction, the value $y_{t+1}$ is revealed and the dataset $\mathcal{D}_{t+1}\coloneq\mathcal{D}\cup\{\mathbf{x}_{t+1},y_{t+1}\}$ is updated. In this case, we can compute $w_{k}^{\text{BMA}}$ recursively using posterior predictive distributions, which allows us to apply exact BMA at time $t$ using the currently-available information: denoting the weights at time $t$ as $w_{t,k}$, with $w_{0,k}=\Pr(\mathcal{M}_{k})$, (2) becomes
@@
Two main problems arise when using BMA: first, estimation via Equation˜2 or Equation˜3 requires access to predictive distributions, either through the marginal likelihood (i.e., the prior predictive) or the posterior predictive. This is often a surmountable problem, as marginal likelihoods are usually available in conjugate models, and BMA with approximated marginal likelihoods has also performed well empirically (Gómez-Rubio et al., 2020).
Perhaps more importantly, BMA is only optimal in terms of predictive error in the so-called “M-closed” setting, where the data were generated by one of the models $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$ (Yao et al., 2018; Minka, 2000). Indeed, in the limit of infinite data, BMA weights concentrate on the single model that most closely resembles the data-generating process. We provide a novel empirical Bayes-based analysis of this fact in Section˜4. This may result in arbitrarily poor posterior predictive accuracy of an ensemble using $\mathbf{w}^{\text{BMA}}$ relative to some other set of weights $\mathbf{w}^{*}$. Additionally, O-BMA can numerically collapse to the “wrong” model, never recovering due to numerical underflow (Waxman & Djurić, 2024).
@@ 2.2 Bayesian Stacking
Bayesian stacking (Yao et al., 2018) (similarly explored by (Clyde & Iversen, 2013; Le & Clarke, 2017)) presents an alternative way to derive a weight vector $\mathbf{w}$. In particular, Bayesian stacking finds the optimal weight vector $\boldsymbol{\mathbf{w}}^{*}$ in the (K1)(K-1)-dimensional simplex $\mathbb{S}^{K}$ by maximizing some score $\mathcal{S}(\boldsymbol{\mathbf{w}},\mathcal{D}^{\prime})$ over a separate dataset $\mathcal{D}^{\prime}$:
@@
Particular attention is given to the log-score; for the predictive dataset (i.e., a holdout or validation set) $\mathcal{D}^{\prime}=\{(\boldsymbol{\mathbf{x}}_{n},y_{n})\}_{n=1}^{N}$ [^1] We will typically use $n=1,\dots,N$ to denote datasets that are not processed sequentially/online, and $t=1,\dots,T$ for their sequential/online counterparts., the corresponding optimization problem is
@@
where $p_{k}(y_{n}\,|\,\boldsymbol{\mathbf{x}}_{n},\mathcal{D})$ is the predictive distribution of model $\mathcal{M}_{k}$. This can be recognized as finding the mixture of estimators with the highest predictive likelihood over $\mathcal{D}^{\prime}$. This, in turn, minimizes the KL divergence between the mixture $q(y\,|\,\boldsymbol{\mathbf{x}})=\sum_{k}w_{k}p_{k}(y\,|\,\boldsymbol{\mathbf{x}},\mathcal{D})$ and the predictive distribution, $p(y\,|\,\boldsymbol{\mathbf{x}})$, which represents the (unknown) true generating mechanism of $y$ (Yao et al., 2018).
The optimization problem (5) is convex and provides nice asymptotic guarantees, but “wastes” data by requiring two separate datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. Yao et al. (2018) address this by showing the score in (5) is well-approximated by the leave-one-out (LOO) predictive density $p(y_{n}\,|\,\boldsymbol{\mathbf{x}}_{n},\mathcal{D}_{-n})$, where $\mathcal{D}_{-n}=\mathcal{D}\setminus\{(\boldsymbol{\mathbf{x}}_{n},y_{n})\}$. For many models, the LOO predictive density can be efficiently estimated with Pareto-smoothed importance sampling (Vehtari et al., 2024), resulting in an efficient optimization problem with a single dataset $\mathcal{D}$. The resulting mixture of estimators has empirically shown performance superior to that of other combination methods, leading to several recent applications.
@@ 2.3 Bayesian Stacking for Time Series & Online Bayesian Stacking
In the published discussion to Yao et al. (2018), Ferreira (2018) discusses Bayesian stacking for time series, where forecasting densities naturally serve as predictive distributions, i.e., (5) becomes
@@
where the first $t^{*}$ data are devoted to (pre-)training each model, and $\mathcal{D}_{1:t-1}$ denotes the first t1t-1 data points. This has notable similarities to pooling methods described by Hall & Mitchell (2007); Geweke & Amisano (2011), with the exception of the summation starting with index $t=1$, reminiscent of the LOO approach. This method was applied by Geweke & Amisano (2012), where optimal weights were computed quarterly. “Windowed” approaches for autocorrelated time series data, where only the last $T_{p}$ points are considered, have also been deployed (Jore et al., 2010; Aastveit et al., 2014), but these are suboptimal in the more general case where data may be exchangeable conditioned on $\boldsymbol{\mathbf{x}}_{t}$. Related to windowed approaches is dynamic model averaging, which is O-BMA with forgetting factors (Raftery et al., 2010). To the best of our knowledge, Bayesian stacking and its time series variants have not been applied in the online setting, where weights are estimated as new data become available. We will refer to approaches to the online problem as online Bayesian stacking (OBS).
As we will see, OBS is a special case of online convex optimization (OCO) (Hazan, 2022), where “learning from experts” is well-studied. We show that the popular Hedge algorithm (Freund & Schapire, 1997) generalizes BMA with a learning rate, but optimizes a different loss function from OBS. Our proposed OBS is differentiated from existing approaches in several ways. First, it directly emulates Bayesian stacking, of recent interest to the Bayesian community. Second, our connection to OCO yields an extremely efficient implementation, unlike methods based on data-driven portfolio selection (e.g., Baştürk et al. (2019)) that rely on particle filtering. Finally, our approach is more general for online learning, whereas data-driven strategies rely on the autocorrelation of time series data.
@@ 3 Insights from Online Convex Optimization
As Bayesian stacking is natively framed as an optimization problem, it is natural to study OBS from the optimization perspective as well. It turns out that our studies here are fruitful: by interpreting posterior predictive values as asset prices, OBS corresponds exactly to a classical problem of portfolio selection. An OCO perspective additionally rediscovers a development of Vovk (2001), which shows that BMA is exactly the Hedge algorithm with a specific choice of learning rate. We additionally discuss a hybrid approach meant to detect whether the model class is M-open or M-closed (or, more precisely, if OBS mixtures will outperform expert selection).
@@ 3.1 The Portfolio Selection Problem
We will first review the basic ideas of online portfolio selection (OPS), but our discussion will necessarily be brief; the interested reader is referred to the recent survey of Li & Hoi (2014), or the excellent lecture notes of Hazan (2022); Orabona (2019).
Problem Statement Consider an investment manager overseeing $K$ stocks, aiming to maximize total future wealth. They create the best constantly rebalanced portfolio (BCRP), i.e., at each time step they re-allocate current wealth such that $w_{k}$ is in stock $k$ (classic portfolio selection assumes no transaction fees). Assets are allocated based on price-relative vectors $\boldsymbol{\mathbf{r}}_{t}$, where $r_{t,k}$ is stock $k$’s relative value increase from day t1t-1 to $t$. The goal is to maximize the (multiplicative) wealth at time $T$, or more conveniently, the (additive) log-wealth:
@@
In the online version of the portfolio selection problem, we seek to find $\boldsymbol{\mathbf{w}}_{t}$ that minimizes the regret with respect to the optimal weights $\boldsymbol{\mathbf{w}}^{*}$,
@@
Although many approaches explicitly model returns as stochastic processes, others allow $r_{t,k}$ to be chosen arbitrarily, and indeed by an adversary. Algorithms that achieve sublinear regret with respect to any non-negative sequence of returns $r_{t,k}$ are known as universal (Orabona, 2019, p. 144). We draw an analogy between the price-relative vectors $\boldsymbol{\mathbf{r}}_{t}$ and the predictive likelihoods $p(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})$, which make the adversarial case of particular interest to us.
Algorithms The portfolio selection problem has an attractive structure from the optimization perspective: the log wealth is a concave function defined over the simplex, a convex set. OPS therefore falls under the (online) convex optimization (OCO) umbrella (Hazan, 2022; Orabona, 2019). To keep with standard terminology, we will introduce each method as minimizing the convex loss function given by $\log\text{wealth}_{T}$, which is equivalent to maximizing the concave function $\log\text{wealth}_{T}$.
The simplest OCO algorithm is online (sub)gradient descent (OGD), a straightforward extension of the classical gradient descent algorithm to a sequence of losses $\ell_{1},\dots,\ell_{T}$, each a function of some generic quantity $\boldsymbol{\mathbf{\theta}}$ belonging to a convex set $\mathcal{K}$. In our case, $\boldsymbol{\mathbf{\theta}}_{t}$ represents weights $\boldsymbol{\mathbf{w}}_{t}$, $\ell_{t}$ is the negative log-return, $\log(\boldsymbol{\mathbf{w}}_{t}^{\top}\boldsymbol{\mathbf{r}}_{t})$, and $\mathcal{K}$ is the simplex $\mathbb{S}^{K}$. At each time instant, $\boldsymbol{\mathbf{\theta}}_{t}$ is updated in the direction of the gradient $\nabla\ell_{t}$ and then projected back onto the convex set $\mathcal{K}$ (e.g., (Hazan, 2022, Algorithm 2.3)).
In many problems, it turns out to be beneficial to consider regularizers that account for the geometry of $\mathcal{K}$ explicitly, resulting in different update steps. This is an online form of the classical mirror descent (Nemirovsky et al., 1983). When $\mathcal{K}$ is the simplex $\mathbb{S}^{K}$ (as in OPS), one popular choice is entropic regularization (e.g., (Hazan, 2022, Section 5.4.2)). The resulting algorithm is known as Exponentiated Gradients (EG) (Helmbold et al., 1998), which uses the update step
@@
where $\odot$ is the elementwise product, and $\eta$ is a learning rate. For the cost function $\ell_{t}(\boldsymbol{\mathbf{w}})=-\log\boldsymbol{\mathbf{w}}_{t}^{\top}\boldsymbol{\mathbf{r}}_{t}$, we have $\nabla_{\boldsymbol{\mathbf{w}}}\ell_{t}=-\boldsymbol{\mathbf{r}}_{t}/\boldsymbol{\mathbf{w}}_{t}^{\top}\boldsymbol{\mathbf{r}}_{t}$, and therefore,
@@
The EG algorithm is conceptually simple, with relatively good regret bounds whenever the gradient is bounded. However, more modern algorithms, such as the online Newton step (ONS) (Hazan et al., 2007), can provide tighter bounds on the regret.